### v2

# A Comprehensive Evaluation on Quantization Techniques for Large Language Models

Abstract
For large language models (LLMs), post-training quantization (PTQ) can significantly reduce memory footprint and computational overhead. Model quantization is a rapidly evolving research field. Though many papers have reported breakthrough performance, they may not conduct experiments on the same ground since one quantization method solution usually contains multiple components. In addition, it is also important to analyze the connections among numerous existing methods to gain in-depth understanding of this field. To bridge these gaps, in this paper, we timely conduct an extensive review of the state-of-the-art methods and also perform comprehensive evaluations on the same ground to ensure fair comparisons of these methods. To our knowledge, this fair and extensive investigation remains critically important yet significantly underexplored in the research community.
To better understand the connections, first, we decouple the published quantization methods into two steps: pre-quantization transformation and quantization error mitigation. We define the former as a preprocessing step applied before quantization to reduce the impact of outliers, making the data distribution flatter and more suitable for quantization. Quantization error mitigation, on the other hand, involves techniques that offset the errors introduced during quantization, thereby enhancing model performance. Second, we extensively evaluate and analyze the impact of different settings of quantization methods, including granularity and symmetry. Third, we also timely analyze and evaluate the latest MXFP4 and NVFP4 data format and its performance.
Our experimental results first demonstrate that the optimized rotation and scaling yield the best performance for pre-quantization transformation, and combining low-rank compensation with GPTQ occasionally outperforms using GPTQ alone for quantization error mitigation. Second, as the granularity becomes finer, the performance improves, but this also incurs additional storage overhead. We also verified that applying asymmetric quantization to weights yields much smaller benefits compared to applying it to activations, which is consistent with the common practice of using asymmetric quantization primarily to activations. Third, we explore the potential of the latest FP4 quantization and find that the format and precision of scaling factors have a significant impact on the performance. We also show the performance of the optimal pre-quantization transformation strategy for INT4 on FP4, revealing that rotation-based methods demonstrate little improvement in performance for MXFP4 and NVFP4, inspiring further investigation in this field.
keywords:
Model Quantization, Post-training Quantization , LLM
1 Introduction
Large language models (LLMs) have achieved remarkable performance on diverse tasks and are increasingly being applied in various scenarios. Nonetheless, LLMs inference process incurs significant memory and computational overhead, leading to high inference latency and posing challenges for computations on source-limited GPUs.
Post-training quantization (PTQ) is a promising solution to reduce the memory and computational cost of LLMs, enhancing their inference efficiency. LLMs quantization has been studied in a wide spectrum of precision configurations. Representative settings include W4A4 [1, 2, 3, 4], W8A8 [5, 6, 7] and W4A16 [8, 9, 10, 11]: W4A4 quantizes both weights and activations to 4-bit precision (INT4); W8A8 and W4A16 quantize weights to INT8 and INT4 and activations to INT8 and FP16/BF16 respectively. Among these three settings, W4A4 is currently the main focus of research, as it holds great potential for improving memory efficiency and inference speed. However, applying lossless W4A4 quantization to LLMs remains a challenging problem. Thus, we concentrate our efforts on exploring W4A4 quantization.
In this paper, we mainly investigate three dimensions. First, we find that existing methods are often combinations of multiple techniques, since using a single technique alone typically fails to achieve excellent quantization performance. OSTQuant [8] and FlatQuant [12] employ multiple techniques to adjust the weights and activations, including rotation and scaling, and also apply the GPTQ [8], a gradient-based method for mitigating quantization error, to quantize the transformed weights. Besides, many techniques share similar transformations, for instance, Quip# [10] and QuaRot [13] both use rotation, also, AWQ [11] and SmoothQuant [6] both use scaling. To achieve a better understanding of existing methods and make our evaluation on the same ground, we decompose the quantization process into two key steps: pre-quantization transformation and quantization error mitigation. The pre-quantization transformation step refers to preprocessing the data before quantization in order to mitigate the effect of outliers, making the data distribution smoother and more quantization-friendly. The quantization error mitigation step involves applying techniques to compensate for the errors introduced by quantization, thereby reducing the overall quantization error and improving performance. For each step, we select several state-of-the-art techniques and conduct an evaluation of the effectiveness of their combinations.
Second, quantization involves a variety of configurations, such as granularity and whether symmetric or not. Different settings can lead to different performance outcomes and often involve trade-offs between efficiency and accuracy. For different precision levels, the optimal quantization settings may vary. And existing solutions usually do not use exactly the same setting, leading to unfair comparison. To bridge this gap, we conduct a comprehensive evaluation on the same quantization configurations and provide recommendations on selecting appropriate configurations. And we integrate the methods used in this paper into our code library to facilitate performance evaluation.
Third, NVIDIA’s GeForce RTX 50 Series, which is newly released and based on the Blackwell architecture, includes support for the FP4 data format (e.g., NVFP4 and MXFP4) in its Tensor Cores, showing significant advantages over INT4. However, the research on FP4 has not been well explored. To bridge this gap, we conduct research on weight-activation FP4 quantization using the popular methods in INT4, showing the great timeliness of our research. For FP4 quantization, we also observe that the format and precision of scaling factors have a significant impact on the performance.
2 Review on Existing Quantization Methods
In this section, we first review recent quantization algorithms in Section 2.1 in terms of two steps: pre-quantization transformation and quantization error mitigation. Table 1 shows our decomposition of the popular quantization methods. Section 2.2 focuses on settings, including symmetry and granularity of quantization methods. Meanwhile, Table 2 summarizes the settings of existing methods, showing that they are not compared on the same ground. Section 2.3 discusses FP4 quantization and its variants, including MXFP4 and NVFP4.
Methods Pre-quantization Transformation Quantization Error Mitigation
SmoothQuant [6] Scaling RTN
Outlier Suppression+ [7] Shifting + Scaling RTN
GPTQ [8] - GPTQ
ZeroQuant-v2 [14] - Low-rank
QuIP [9] Rotation GPTQ
AWQ [11] Scaling RTN
Atom [1] Reorder GPTQ
OmniQuant [2] Shifting+Scaling GPTQ
Qserve [15] Scaling + Rotation GPTQ
LQER [16] Scaling Low-rank
QuaRot [13] Rotation GPTQ
SpinQuant [3] Rotation GPTQ
ResQ [17] Reorder + Rotation GPTQ
FlatQuant [12] Scaling + Rotation GPTQ
OSTQuant [4] Scaling + Rotation GPTQ
Table 1: Two steps decomposition of quantization methods.
2.1 Two-steps of quantization
2.1.1 Pre-quantization transformation
Pre-quantization transformation is a step before quantization, applying operations such as shifting, scaling, and rotation to process weights and activations. These transformations aim to flatten and smooth the hard-to-quantize tensors, thereby enhancing the quantization performance. LLMs are difficult to quantize due to the outliers in the activations. Compared to most of the activation values, the magnitude of outliers is substantially higher, often exceeding 100x larger in some layers. The outliers will dominate the maximum magnitude measurement, leading to low effective quantization levels. This pre-quantization transformation step is illustrated in Fig. 1.
Methods Precision Symmetric Granularity
W A
SmoothQuant [6] W8A8 T F T,C
GPTQ [8] W3A16,W4A16 T - G
ZeroQuant-v2 [14] W4A16,W3A16,W2A16,W4A8 T,F T,F G
QuIP [9] W2A16,W3A16,W4A16 T - G
AWQ [11] W3A16,W4A16 T - G
Atom [1] W3A3,W4A4 T F G
OmniQuant [2] W2A16,W3A16,W4A16,W4A4,W6A6,W8A8 F F C
Qserve [15] W4A8 T F C
LQER [16] W4A6,W4A8 T F G
QuaRot [13] W4A4 T F C,G
SpinQuant [3] W4A4,W4A8 T F C
ResQ [17] W4A4 T F C
FlatQuant [12] W4A4 T F C
OSTQuant [4] W4A4 T F C
Table 2: Symmetry and granularity of quantization methods. In granularity, T stands for per-tensor, C represents per-channel or per-token, and G means per-group.
Figure 1: Pre-quantization transformation: a process that transforms activations and weights to make them easier to quantize. The left graph illustrates a tensor that is difficult to quantize due to the presence of numerous outliers, while the right graph shows that after the pre-quantization transformation, the tensor becomes more uniformly distributed, making quantization easier.
Shifting
Activation distribution is often asymmetric, and this asymmetry varies across different channels. Aligning the center of each channel can improve the performance of quantization, since we quantize activations per-token rather than along the channel dimension. Outlier Suppression+ [7] proposes shifting activation values to eliminate the asymmetry feature across channels. The shifting process is to use a vector $T$ that transfers the channels of the activation. For a linear layer, it is formulated as:
[…]
where the layer input $X\in\mathbb{R}^{T\times C_{in}}$ is shifting and becomes $\hat{X}$. $T\in\mathbb{R}^{1\times C_{in}}$ is a vector representing the channel-wise shifting parameters. To keep mathematical equivalence, an additional term, the product of $T$ and $W\in\mathbb{R}^{C_{in}\times C_{out}}$, is added to the bias $B\in\mathbb{R}^{1\times C_{out}}$.
Methods Shifting Strategy Shifting parameters $T$
OS+ [7] Pre-scaling shifting $T_{i}=\frac{\min(X_{i})+\max(X_{i})}{2}$
OmniQuant [2] Pre-scaling shifting Gradient-based optimization
Table 3: Comparison of shifting-based quantization methods.
As shown in Table 3, Outlier Suppression+ [7] uses calibration to get the vector $T$, while OmniQuant [2] employs learnable channel-wise shifting for outlier suppression, combined with the learnable scaling technique discussed in the next paragraph, and jointly optimizes these two kinds of pre-quantization transformation parameters.
Scaling
Scaling is a commonly used pre-quantization transformation method, which scales activations in different input channels before quantization, and inversely scales the corresponding weight channels to make the results stay unchanged. SmoothQuant [6] first proposes using scaling to smooth the quantization difficulty from activations to weights. Compared to weights, activations have more outliers that make quantization harder. It is observed that outliers in activations appear in certain channels, so applying channel-wise scaling to activations can eliminate the outliers to a certain degree, using the following equation for linear layers:
$Y=XW=(\underbrace{X\cdot S^{-1}}_{\hat{X}})(\underbrace{S\cdot W}_{\hat{W}})=\hat{X}\hat{W}$ (2)
where $S$ is a diagonal matrix composed of scaling factors of different channels, which can be obtained in various ways.
Methods Scaling Strategy Scaling parameters $S$
SmoothQuant [6] Scaling only $\begin{aligned} &S_{i}=\frac{\max(|X_{i}|)^{\alpha}}{\max(|W_{i}|)^{1-\alpha}}\end{aligned}$
AWQ [11] Scaling only $\begin{aligned} &S_{i}=|X_{i}|_{avg}^{\alpha}\\ &\alpha^{*}=\arg\min_{\alpha}\mathcal{L}(S_{Xi}^{\alpha})\end{aligned}$
OS+ [7] Post-shifting scaling $\begin{aligned} &S_{i}=\max\left(1.0,\frac{\max(X_{i}-T_{i})}{t}\right)\\ &t^{*}=\arg\min_{t}\mathcal{L}(S_{i})\end{aligned}$
OmniQuant [2] Post-shifting scaling Gradient-based optimization
FlatQuant [12] Pre-rotation scaling Calibration initialization and gradient-based optimization
OSTQuant [4] Post-rotation scaling Gradient-based optimization
Table 4: Comparison of scaling-based quantization methods.
As shown in Table 4, SmoothQuant [6] uses calibration to compute $S$. AWQ [11], as a weight-only quantization, approaches the problem from the perspective of protecting salient weights. Its scaling parameter is determined by the average magnitude of the activation, while the optimal exponent $\alpha^{*}$ is found by minimizing the loss through grid searching. Outlier Suppression+ [7] also uses grid searching optimization to find the optimal scaling after applying shifting. Unlike adopting the hand-crafting parameters in calibration, Outlier Suppression+ [7] uses grid searching optimization to find the optimal scaling after applying shifting. OmniQuant [2] optimizes the quantization parameters via gradient-based methods to minimize the block-wise quantization error. FlatQuant [12] and OSTQuant [4] both use learnable per-channel scaling, but at different stages of pre-quantization transformation, leading to different practical meanings.
Rotation
The rotation operation is a technique that adjusts data distribution by multiplying the data matrix by an orthogonal matrix. This operation largely changes the activation distributions, reduces outliers in the data, and thereby enhances quantization accuracy. The rotation process in a linear layer can be formulated as:
$Y=XW=(\underbrace{XO}_{\hat{X}})(\underbrace{O^{\top}W}_{\hat{W}})=\hat{X}\hat{W}$ (3)
where the input $X$ is rotated with an orthogonal matrix $O$, resulting in $\hat{X}=XO$. To preserve the model’s output, the weight $W$ is simultaneously transformed by $O^{\top}$, which is the transpose of $O$.
Methods Rotation Strategy Rotation parameters $O$
QuIP [9] Rotation only Random orthogonal matrices
QuIP# [10] Rotation only Random Hadamard matrices
QuaRot [13] Rotation only Random Hadamard matrices
SpinQuant [3] Rotation only Gradient-based optimization
FlatQuant [12] Post-scaling Rotation Gradient-based optimization
OSTQuant [4] Pre-scaling Rotation Gradient-based optimization
Table 5: Comparison of rotation-based quantization methods.
As shown in Table 5, QuIP [9] first introduced rotation with random orthogonal matrices to reduce incoherence in weight-only quantization. QuIP# [10] replaces this with random Hadamard matrices for greater efficiency. QuaRot [13] extends rotation to both weights and activations quantization using random Hadamard matrices, effectively eliminating outliers. SpinQuant [3] further improves stability by optimizing the rotation matrix on the Stiefel manifold to minimize quantization-induced performance variance, while still defaulting to random Hadamard for a few online rotations.
Rotation is a highly effective pre-quantization transformation technique for reducing outliers in low-bit quantization, and it has also been widely adopted in many other quantization methods together with other pre-quantization transformation techniques to further improve performance (e.g., ResQ [17], QServe [15], FlatQuant [12], and OSTQuant [4], etc).
2.1.2 Quantization error mitigation
The quantization error mitigation step employs techniques that mitigate the error introduced by quantization by compensating for it. This error between the original model and the quantized model can be computed from calibration datasets, and compensated to the original weights in linear layers or through low-rank branches. Fig. 2 demonstrates two commonly used quantization error mitigation methods.
Figure 2: Quantization error mitigation: compensate the error produced by quantization. Here, $dq(W_{q})$ denotes dequantizing $W_{q}$ back to the original precision value space, for computational purposes.
Self-compensation
We define self-compensation as directly compensating for errors on the original weights. Existing methods that employ self-compensation are generally based on analyzing the loss error introduced by quantization and make use of its Hessian-based quadratic approximation, which can be formulated as:
$\Delta\mathcal{L}=\mathcal{L}(dq(W_{q}))-\mathcal{L}(W)\approx\frac{1}{2}(dq(W_{q})-W)^{\top}H(dq(W_{q})-W)$ (4)
where $\mathcal{L}(W)$ is the loss function of the model evaluated at weights $W$. $W_{q}$ and $dq(W_{q})$ denote the quantized weights and their dequantized version, respectively. Self-compensation methods need to find an optimal weight perturbation $\delta$ to add to $W$, such that the quantization loss error in Eq. (4) is minimized. Table 6 summarizes the computation strategies of $\delta$ adopted by various self-compensation quantization methods.
Methods Weight perturbation $\delta$ computation
OBQ [18] Layer-wise MSE
GPTQ [8] Layer-wise MSE
QuIP [9] LDL decomposition
GPTAQ [19] Consider input deviation
GQuant [20] Add gradient information
Table 6: Comparison of self-compensation quantization methods.
Some methods simplify the quantization problem to minimize the squared error between the quantized and original output of each layer [18, 8]. Based on the squared error loss function, OBQ [18] and GPTQ [8] compute the Hessian matrix used in Eq. (4) and further calculate the optimal weight perturbation $\delta$. Unlike OBQ, which uses greedy orders to quantize weights, GPTQ quantizes all rows of weights in a fixed order to achieve a faster quantization process.
Some other methods take into account additional factors beyond the linear layer-wise formulation. While also aiming to minimize the quadratic proxy of the weight error in Eq. (4), QuIP [9] performs an LDL decomposition of the Hessian matrix to construct a compensation term that corrects for quantization-induced errors. GPTAQ [19], built upon GPTQ, addresses the limitation that GPTQ’s layer-wise optimization objective ignores the deviation between a layer’s actual input and that under the full-precision model. GQuant [20] further observes that the layer-wise MSE treats all hidden features equally, and therefore incorporates per-feature gradients into the objective function to improve quantization accuracy.
Low-rank decomposition compensation
Beyond compensating errors directly on the original weights, an alternative approach is to introduce an auxiliary branch dedicated to error correction.
Some prior works leverage a low-rank term to reconstruct quantization error (e.g., ZeroQuant-v2 [14], CALDERA [21], LQER [16], QERA [22], etc), where the low-rank branch can either be quantized to higher precision, such as INT8, or remain in FP16. After incorporating a low-rank branch, the output $Y$ is computed as:
$Y=Xdq(W_{q})+XAB$ (5)
where $A\in\mathbb{R}^{C_{in}\times r}$, $B\in\mathbb{R}^{r\times C_{out}}$ are low-rank matrices, and $r\ll\min(d,k)$ is the bottleneck rank. The product $AB\in\mathbb{R}^{d\times k}$ serves as a low-rank approximation of the quantization error. Table 7 illustrates the computation process of the low-rank terms in some quantization methods.
Methods Low rank terms $A,B$ computation
ZeroQuant-v2 [14] $\text{SVD}(dq(W_{q})-W)$
LQER [16] $\text{SVD}(dq(W_{q})-W)$
L2QER [16] $\text{SVD}(S(dq(W_{q})-W)),A:=S^{-1}A$
CALDERA [21] LPLR factorization [21]
QERA [22] $\text{SVD}(R^{\frac{1}{2}}_{X\!X}(dq(W_{q})-W)),A:=(R^{\frac{1}{2}}_{X\!X})^{-1}A$
Table 7: Comparison of low-rank decomposition compensation quantization methods. $A,B$ represents the low rank terms. $R_{X\!X}$ denotes the autocorrelation matrix of the input $X$.
ZeroQuant-v2 [14] and LQER [16] apply SVD composition to reconstruct the quantization error matrix $E=W-W_{q}$. Based on LQER, L2QER, which is proposed in LQER, scales the $E$ before applying SVD heuristically, and inversely applies scales in $A_{k}$. This operation aims to more precisely approximate the quantization error corresponding to the salient weights.
In addition to the aforementioned approaches, CALDERA [21] and QERA [22] derive the optimal low-rank terms for quantization error mitigation by framing it as an optimization problem that minimizes the squared error between each layer’s output in Eq. (5) and its original output.
In contrast to approaches that introduce low-rank branches externally, QwT [23] integrates lightweight linear layers within each decoder layer block of LLMs, directly compensating quantization errors, offering new insights for quantization error mitigation.
2.2 Settings of quantization
2.2.1 Symmetric vs. asymmetric
Uniform quantizers can be broadly categorized into symmetric and asymmetric schemes, depending on how the zero point is handled during the mapping from floating-point to integer values.
(a) Symmetric quantization
(b) Asymmetric quantization
Figure 3: Symmetry of quantization
Symmetric quantization forces the zero point to align with the floating point, resulting in a quantization range that is symmetric with respect to zero. Asymmetric quantization allows the zero point to shift dynamically according to the minimum and maximum values of the data. This enables a tighter fit to the data distribution, especially when it is skewed, at the cost of slightly more complex computation.
Symmetric quantization enables a very simple dequantization process. For linear layers, dequantization can be efficiently performed by multiplying the quantized output by the product of the scale factors for the weights and activations This simplifies computation and is often preferred in hardware implementations. However, symmetric quantization may lead to suboptimal representation when the data distribution is not centered around zero.
Asymmetric quantization incurs additional computational cost during dequantization, while it often outperforms symmetric quantization in terms of accuracy. As such, the selection of quantization symmetry should be made with consideration of the specific use case and performance-complexity trade-offs.
2.2.2 Granularity of quantization
The granularity of quantization often largely affects the low-bit quantization performance. Commonly adopted granularities include per-tensor, per-channel, or per-token, and per-group, as illustrated in Fig. 4.
Figure 4: Granularity of quantization: (1) Per-tensor quantization; (2) Per-channel quantization; (3) Per-group quantization.
In per-tensor quantization, a single set of affine transformation parameters is applied to the entire tensor, with the scale ($\Delta$) and zero point ($Z$) computed from the tensor’s minimum and maximum values. A finer-grained alternative is per-channel quantization for weights and per-token quantization for activations, where each input channel of the weights or each token of the activations shares the same quantization parameters. For low-bit quantization scenarios, even finer granularity, per-group quantization, is often employed to further enhance performance. In this case, the elements within a single channel or token are subdivided into groups of a predefined size, with each group quantized independently using its own parameters.
Finer granularity typically leads to improved performance but also incurs higher memory overhead due to the need to store more quantization parameters. Besides, as QServe [15] mentioned, per-group quantization will also raise computational cost at the hardware level. Therefore, while finer granularity can enhance accuracy, it may also reduce the overall efficiency of the system.
In weight-only quantization, papers such as GPTQ [8], AWQ [11], and QuIP [9] all employ per-group quantization to further enhance performance. For weight-activation quantization, Atom [1] also adopts a per-group quantization strategy. However, with the introduction of rotation-based methods in QuaRot [13] and SpinQuant [3], the performance of per-channel quantization has been significantly improved. Inspired by this, subsequent papers like FlatQuant [12], ResQ [17], and OSTQuant [4] have also adopted per-channel granularity for quantization.
2.3 FP4 Quantization and Variant Formats
Recently, NVIDIA’s GeForce RTX 50 Series based on NVIDIA’s Blackwell architecture [24] has been released, which has enabled very low-precision operations using FP4 variants like MXFP4 [25, 26] and NVFP4 [27]. Both MXFP4 and NVFP4 store each value in an E2M1 format, which consists of one sign bit, one mantissa bit, and two exponent bits. Unlike the INT4 format, they adopt a fixed granularity and use symmetric per-group quantization, with MXFP4 employing a group size of 32 and NVFP4 using a finer granularity of 16. Each group is associated with a scaling factor, and this rescaling process is hardware-accelerated. MXFP4 uses an FP8 (E8M0) format, while NVFP4 uses an FP8 (E4M3) format. Additionally, NVFP4 incorporates an extra FP32 scaling factor per tensor to extend its representable range.
The value of an encoding FP4 (E2M1) data is represented as:
$\text{value}=\begin{cases}(-1)^{S}\times 2^{1-\text{bias}}\times(0+2^{-m}\times M)&\text{if }E=0\\ (-1)^{S}\times 2^{E-\text{bias}}\times(1+2^{-m}\times M)&\text{if }E>0\end{cases}$ (6)
where $S,E,M$ correspond to the sign, exponent, and mantissa fields, respectively. The bias is the exponent bias, which is set to 1. So the FP4 (E2M1) encoding values are [-6, -4, -3, -2, -1.5, -1, -0.5, 0, 0.5, 1, 1.5, 2, 3, 4, 6].
FP4 can better handle long-tail distributions compared to INT4, as its values are non-uniformly distributed. In LLMs, both weights and activations approximately follow normal distributions, with activations particularly exhibiting long-tail characteristics. Fig. 5 shows the weight distribution under BF16, INT4, and MXFP4. A clear observation is that under INT4 quantization, values with larger absolute magnitudes represent fewer data points, whereas MXFP4 more fully utilizes the representational capacity of 4-bit precision. Therefore, FP4 is more suitable than INT4 as a quantization data type.
(a) Full-precision weight
(b) INT4 quant weight
(c) MXFP4 quant weight
Figure 5: Weight distribution of a q_proj layer of LlaMA-3.2-1B. (a): Weight distribution in full-precision(BF16). (b): Weight distribution after INT4 quantization (group size=32), compared with MXFP4. (c): Weight distribution after MXFP4 quantization.
LLM-FP4 [28] studies FP4 precision quantization for LLMs, including searching for the optimal exponent bias and maximal quantization value. It also proposed a pre-shifted exponent bias technique to address the issue of high inter-channel variance. FP4-Quantization [29] improves the FP4 data representation while introducing a method called LREC (low-rank error calibration) to quantize weights into FP4.
3 Extensive Evaluations
Quantization settings We conduct experiments using W4A4 precision, comparing the results with the original precision (BF16 or FP16). In the main results, we use per-token or per-group (group size=128) asymmetric quantization for activations. For weight quantization, we use per-channel or per-group (group size=128) symmetric quantization, where we extract the clipping ratio using a linear search over the squared error.
Models and tasks. We conduct experiments on LLaMA-3.2-1B, LLaMA-3.2-3B, and Qwen-2.5-3B. We evaluate their perplexity on the WikiText-2 dataset [30] and assess their accuracy on several zero-shot evaluation tasks, including commonsense reasoning, mathematics, sentiment analysis, and question answering. Specifically, we utilize the following datasets for evaluation: MMLU [31], SST2 [32], MathQA [33], ARC‑C [34], HellaSwag [35], WinoGrande [36], BoolQ [37], LAMBADA [38], and PIQA [39].
Implementation. To calibrate the GPTQ [8] quantization and the low-rank compensation branch, we use 128 samples from WikiText-2 with a sequence length of 2048. The training of the learnable orthogonal and scaling matrices follows the OSTQuant [4] procedure, using 1,000 samples from WikiText-2 and optimized with RiemannAdam. For the low-rank decomposition, we set the rank k=32 and store the low-rank matrices in FP16. For the non-trainable scaling matrix, we follow the SmoothQuant [6] calibration scheme with $\alpha$=0.5.
3.1 Overall performance
Evaluate Methods Channel Group-128
PT QEM Wiki2‑PPL($\downarrow$) 0‑shot9 Avg.($\uparrow$) Wiki2‑PPL($\downarrow$) 0‑shot9 Avg.($\uparrow$)
BF16 9.76 54.78 9.76 54.78
Rotation RTN 17.83 45.24 14.84 47.51
Rotation Low-rank 17.16 45.47 14.5 47.04
Rotation GPTQ 12.95 47.79 11.65 49.38
Rotation GPTQ+Low-rank 12.81 48.35 11.61 50.49
Rotation+Scaling RTN 14.88 46.41 13.3 49.22
Rotation+Scaling Low-rank 14.58 46.76 13.07 49.95
Rotation+Scaling GPTQ 11.8 48.89 10.89 50.63
Rotation+Scaling GPTQ+Low-rank 11.73 49.06 10.87 51.83
Table 8: Evaluation of Quantization Method Combinations, in which rotation or scaling is optimized, in terms of Perplexity on WikiText2 and averaged accuracy on nine Zero-Shot tasks across per-channel and per-group settings of LlaMA-3.2-1B. PT and QEM represent pre-quantization transformation and quantization error mitigation, respectively.
Evaluate Methods Channel Group-128
PT QEM Wiki2‑PPL($\downarrow$) 0‑shot9 Avg.($\uparrow$) Wiki2‑PPL($\downarrow$) 0‑shot9 Avg.($\uparrow$)
BF16 9.76 54.78 9.76 54.78
Rotation RTN 20.31 43.26 15.18 46.62
Rotation Low‑rank 19.84 43.09 14.83 46.87
Rotation GPTQ 13.31 48.02 11.96 50.05
Rotation GPTQ+Low‑rank 13.28 48.37 11.98 50.48
Rotation+Scaling RTN 20.96 42.87 16.25 45.01
Rotation+Scaling Low‑rank 21.97 43.07 15.55 45.46
Rotation+Scaling GPTQ 14.04 47.53 12.23 48.95
Rotation+Scaling GPTQ+Low‑rank 13.91 46.46 12.17 49.41
Table 9: Evaluation of Quantization Method Combinations, in which rotation or scaling is non-optimized, in terms of Perplexity on WikiText2 and averaged accuracy on nine Zero-Shot tasks across per-channel and per-group settings of LlaMA-3.2-1B. PT and QEM represent pre-quantization transformation and quantization error mitigation, respectively.
We select two of the most widely used pre-quantization transformation techniques, scaling and rotation, and combine them with two quantization error mitigation methods: GPTQ and low-rank decomposition compensation. A series of experiments is conducted to evaluate these combinations. Table 8 presents the results of using optimized scaling and rotation, illustrating the performance of different method combinations. Table 9 presents the results obtained with randomly initialized rotations and calibrated scaling factors.
Our experiments reveal that the best performance is achieved when both optimized scaling and rotation are applied in conjunction with GPTQ and the low-rank decomposition. This configuration yields a perplexity of 11.73 in per-channel quantization for the Llama-3.2-1B model, which is close to the BF16 baseline of 9.76.
Between the two quantization error mitigation methods, GPTQ clearly plays a more significant role than low-rank adaptation. Under the optimal pre-quantization transformation setting (optimized rotation and scaling), using GPTQ alone results in perplexity of 11.8, whereas using only the low-rank method yields 14.58, under the per-channel settings. When both methods are used together, performance improves further, from 11.8 to 11.73 in the per-channel setting, and from 10.89 to 10.87 in the per-group setting, narrowing the gap to the full-precision model.
In addition, using both optimized rotation and scaling improves performance over using only optimized rotation. However, when employing non-optimized rotation and scaling, rotation alone tends to yield better results than the combination, which indicates that calibrated scaling may not perform well together with rotation in W4A4 settings.
Model precision
Wiki2
PPL($\downarrow$)
Zero-Shot Accuracy($\uparrow$)
General Sentiment Math Reasoning QA Avg.
MMLU SST2 MathQA PiQA HellaS WinoG BoolQ Lambada Arc-C
Llama-
3.2-1B
fp16 9.76 36.76 66.86 28.94 74.65 63.67 60.38 63.67 62.1 36.01 54.78
w4a4 11.73 29.21 55.28 25.43 69.15 57.93 56.43 59.02 56.78 32.34 49.06
Llama-
3.2-3B
fp16 7.82 54.06 73.62 34.47 77.48 73.58 69.46 72.97 70.08 46.08 63.53
w4a4 8.9 46.7 68.58 30.55 73.99 69.54 65.43 68.75 65.32 41.55 58.93
Qwen-
2.5-3B
fp16 8.03 65.13 90.37 37.15 78.73 73.55 68.75 77.58 66.8 47.35 67.27
w4a4 9.22 56.32 69.15 32.5 74.16 67.52 62.12 65.41 57.77 42.83 58.64
Table 10: Comparison of symmetric and asymmetric quantization in Terms of Perplexity on WikiText2 and averaged accuracy on Zero-Shot tasks of LlaMA-3.2-1B
In Table 10, we further conduct experiments on different models: LLaMA-3.2-1B, LLaMA-3.2-3B, and Qwen-2.5-3B, using the optimized scaling and rotation, and combined them with GPTQ and an additional low-rank decomposition.
3.2 Symmetric vs. asymmetric
Granularity Metric rotation+scaling+GPTQ Rotation+Scaling+GPTQ+Low-rank
Sym W-Asym A-Asym Asym Sym W-Asym A-Asym Asym
channel Wiki2-PPL($\downarrow$) 12.73 12.7 11.8 11.7 12.63 12.65 11.73 11.64
Arc-Challenge Score(%) 31 31.7 33.3 33.7 31.9 29.9 32.3 32.7
Table 11: Comparison of symmetric and asymmetric quantization in terms of Perplexity on WikiText2 and Arc [34] accuracy of LlaMA-3.2-1B
In Table 11, we evaluate the performance of applying symmetric or asymmetric quantization to weights and activations, respectively. The results demonstrate that asymmetric quantization for activations leads to a significant improvement over symmetric quantization, while the performance gain for weights is relatively minor. We attribute this to the fact that activation distributions tend to be asymmetric, whereas weights are generally more symmetric in distribution.
Therefore, an effective configuration is to use symmetric quantization for weights and asymmetric quantization for activations. This setup not only achieves much better performance than fully symmetric quantization, but also reduces the computational overhead associated with the dequantization process, compared to fully asymmetric quantization for both weights and activations.
3.3 Granularity
As illustrated in Fig. 6, we conduct a comprehensive comparison of per-group quantization performance across varying group sizes, measured by perplexity on WikiText2. The experimental results reveal that finer quantization granularity yields superior model performance (lower perplexity), albeit at the expense of elevated storage requirements for quantization parameters. This trade-off stems primarily from the scaling factors, whose storage overhead scales inversely with group size. Quantitative analysis of these additional bit costs is presented in Table 12, which compares the average bit overhead introduced by scaling factors under symmetric weight quantization.
Figure 6: Comparison of different group sizes of quantization in terms of Perplexity on WikiText2 and averaged accuracy on Zero-Shot tasks of LlaMA-3.2-1B
Granularity group-32 group-64 group-128 group-256 group-512
Average Extra bits
1 0.5 0.25 0.125 0.0625
Table 12: Average extra bits introduced by FP32 scaling factors of different granularities
3.4 FP4 vs. INT4
Extending our experiments to the FP4 format, we compare several variants of FP4, including MXFP4, NVFP4, and naive FP4 implementation with varying granularities and scaling factor formats. For per-channel quantization, we employ asymmetric quantization for activations and symmetric quantization for weights. To ensure a fair comparison with MXFP4 and NVFP4, all other experiments using per-group quantization adopt symmetric quantization for both activations and weights.
First, we apply no extra pre-quantization transformation, conducting experiments using only quantization error mitigation methods. Table 13 demonstrates that FP4 (per-channel) shows much better performance than INT4 (per-channel). We believe this is because, in the absence of the pre-quantization transformation, FP4 better aligns with the data distribution, thus achieving superior results. For finer granularity scenarios, FP4 and INT4 show very small differences, and both the format of the scaling factor and the size of the group significantly impact performance. Compared to MXFP4, NVFP4 achieves better performance due to its smaller group size and more precise scaling factor format. With the E4M3 variant of FP8, NVFP4 can encode scaling factors with fractional precision. By contrast, MXFP4 uses the E8M0 format, which restricts scale factors to powers of two, resulting in much coarser quantization. Moreover, NVFP4 employs a two-level micro-block scaling strategy: it first applies a per-tensor scaling factor (FP32) to compensate for the smaller range that E4M3 scales can represent. NVFP4 performs better than MXFP4, showing a lower risk of noticeable accuracy drop, matching NVIDIA’s official conclusions [40].
Quantization
error
mitigation
Data Format
INT4
channel, FP32
FP4
channel, FP32
INT4
g16, FP32
FP4
g16, FP32
MXFP4
g32, E8M0
FP4
g32, E4M3
FP4
g16, E8M0
NVFP4
g16, E4M3
RTN 256.97 135.68 11.75 11.34 15.17 12.04 14.52 11.54
GPTQ 161.04 101.36 11.47 10.87 13.24 11.36 12.77 10.98
low-rank 179.87 132.14 11.69 11.32 15.56 11.98 14.67 11.48
GPTQ+low-rank 139.31 96.33 10.92 10.76 12.69 11.18 12.25 10.83
Table 13: Evaluation of INT4 and FP4 format quantization without using pre-quantization transformation, in terms of Perplexity on WikiText2 of LlaMA-3.2-1B. Channel means quantization granularity is per-channel; g16 and g32 denote per-group quantization with group sizes of 16 and 32, respectively. FP32, E8M0, and E4M3 refer to the format of the scaling factor. And for E4M3, the scaling strategy follows NVFP4’s approach, employing a two-level scaling scheme.
Quantization
error
mitigation
Data Format
NVFP4
W:E4M3, A:E4M3
FP4
W:E2M1, A:E4M3
FP4
W:E2M1, A:E2M1
FP4
W:INT4, A:E4M3
FP4
W:INT4, A:INT4
RTN 11.54 14.10 4130 14.66 4777
GPTQ 10.98 13.38 3618 17.96 3803
Low-rank 11.48 13.75 4057 14.43 5356
GPTQ+Low-rank 10.83 11.38 3650 11.59 4124
Table 14: Evaluation of FP4 format quantization using low bit scaling factor and group size 16, in terms of Perplexity on WikiText2 of LlaMA-3.2-1B. E4M3, E2M1, and INT4 refer to the format of the scaling factor. W and A denote weight and activation, respectively. All of these follow NVFP4’s approach, employing a two-level scaling scheme.
We further reduce the bit-width of the scaling factor to 4 bits to evaluate its effectiveness under such a low-bit setting. As shown in Table 14, applying 4-bit scaling factors to both weights and activations leads to very poor performance. When applied only to weights, the performance remains moderate and is worse compared with NVFP4.
Evaluate Methods Wiki2‑PPL($\downarrow$)
Pre-quantization
transformation
Quantization error
mitigation
INT4
channel, FP32
MXFP4
g32, E8M0
NVFP4
g16, E4M3
scaling GPTQ 134.85 13.24 10.97
scaling low-rank 165.4 14.91 11.36
scaling GPTQ+low-rank 124.56 12.58 10.81
optimized rotation+scaling GPTQ 11.8 12.37 11.04
optimized rotation+scaling low-rank 14.58 15.43 12.57
optimized rotation+scaling GPTQ+low-rank 11.73 12.29 11.01
Table 15: Evaluation of INT4, MXFP4, and NVFP4 using both pre-quantization transformation and quantization error mitigation techniques, in terms of Perplexity on WikiText2 of LlaMA-3.2-1B. Channel means quantization granularity is per-channel; g16 and g32 denote per-group quantization with group sizes of 16 and 32, respectively. FP32, E8M0, and E4M3 refer to the format of the scaling factor.
Notably, rotation-based methods yield smaller improvements for MXFP4 and NVFP4 than for INT4 (per-channel). As argued in AMXFP4 [41], grouping activation tensors into small micro-scaled units mitigates outliers like rotation. For MXFP4 and NVFP4 formats, the group size is extremely small, so data rotation contributes little and may even lead to degradation.
4 Conclusions
In this paper, we have presented an experimental evaluation of Large language model quantization. We systematically decouple the commonly used quantization into two key steps: pre-quantization transformation and quantization error mitigation. After a comprehensive review of existing approaches, we choose scaling and rotation in pre-quantization transformation step, GPTQ, and low-rank compensation in the quantization error mitigation step to further conduct experimental evaluations on the same ground. We found that using the optimized rotation and scaling, together with two compensation methods, demonstrates the best performance. For quantization settings, asymmetric quantization and smaller group sizes yield better results. And using asymmetric quantization on activations shows more benefits on performance compared with weights. For FP4 quantization, we can adopt methods used in INT4 quantization. However, rotation-based pre-quantization transformation yields no observable more performance gains for MXFP4 and NVFP4 than scaling-based methods. And we find that the format of scaling factors demonstrates a large effect. For FP8 format scaling factor, FP8 (E4M3) with an additional per-tensor FP32 scaling factor is better than FP8 (E8M0). For small group sizes, using a 4-bit scaling factor for weights is also feasible, albeit with some loss in accuracy. To achieve optimal performance under FP4 quantization, further investigation is required to explore FP4’s full potential and limitations.